### v1

# Hybrid plasmonic nanoantennas using corner reflectors for enhanced single-photon emission

Abstract
The emission rate of atom-like photon sources can be significantly improved by coupling them to plasmonic resonant nanostructures. These arrangements function as nanoantennas, serving the dual purpose of enhancing light-matter interactions and decoupling the emitted photons. However, there is a contradiction between the requirements for these two tasks. A small resonator volume is necessary for maximizing interaction efficiency, while a large antenna mode volume is essential to achieve high emission directivity. In this work, we analyze a hybrid structure composed of a noble metal plasmonic resonant nanoparticle coupled to the atom-like emitter, designed to enhance the emission rate, alongside a corner reflector aimed at optimizing the angular distribution of the emitted photons. A comprehensive numerical study of silver and gold corner reflector nanoantennas, employing the finite-difference time-domain method, is presented. The results demonstrate that a well-designed corner reflector can significantly enhance photon emission directivity while also substantially boosting the emission rate.
1 Introduction
Efficient and reliable single-photon sources, capable of producing on-demand high-purity single-photon states, are key elements for the emerging field of photon-based quantum technologies [1, 2]. Atom-like emitters, based either on solid-state or molecular systems, are a promising alternative for this purpose [3, 4, 5, 6, 7]. In the past years, nanoscale elements have been introduced to enhance the emission rate of the photon sources and to tailor their spatial radiation properties. Such nanoantennas (NAs) reproduce at the nanoscale the action of traditional radiofrequency (RF) aerials. Various traditional antenna geometries have been considered for applications in the optical realm: dipole [8, 9], patch [10], bowtie [11] or the Yagui-Uda array [12, 13, 14].
Antennas for light [12] serve the dual purpose of enhancing the emitter fluorescence and facilitating the out-coupling of the generated photons. The performance requisites for the two functions are antithetical: whereas a small mode volume is required for an enhanced emission rate, high antenna directivity is well known to be the result of a large electrical size, i.e, the extent of the structure measured using the wavelength of the electromagnetic field as unit reference [15, 16]. In [4], such antipodal requirements have been reconciled by photomodifying a nanopatch antenna assembled around a nanodiamond with a nitrogen-vacancy center to reduce the the cavity mode size while keeping a large size emitter to spatially shape the radiation produced. In this work, we tackle this issue by using a hybrid plasmonic structure built from the combination of a conventional NA acting as the feeder element of a corner reflector (CR) [17]. This strategy permits simultaneously optimize the emission rate at the feeder design and to use a CR to enhance the directivity of the photons produced.
The use of CRs in antenna design has been prevalent since Kraus introduced them in 1940 [17], owing to their outstanding performance, simplicity, and ease of construction. This work demonstrates that these advantageous properties can also be applied to nanoscale photonics. The potential of CRs to enhance photon emission was highlighted in [18], where the radiation patterns generated by an electric dipole emitter with a perfect conductor CR were analyzed. Here, we present a thorough numerical study of the proposed NA, utilizing two alternative noble metals: silver and gold. The feeder is assumed to be constituted by an atom-like dipole emitter coupled to a spherical nanoparticle (NP), though the proposed CRs can also accommodate more complex feeder geometries, such as dimers or bowtie structures [8, 11], which may offer enhanced emission. The analysis is conducted using the finite-difference time domain method (FDTD)[19, 20], enabling the calculation of the radiative emission rate enhancement factor, losses within the metallic structures, and the far-field patterns of the emitted photons. This approach provides a comprehensive characterization of the NA.
The geometry employed in the analyses is depicted in Fig. 1. A spherical NP with diameter $d$ is positioned at the origin. The atom-like emitter is represented by a point dipole source $\mathbf{p}$ located at a distance $s$ from the surface of the metal nanosphere that is placed at a distance $S$ from the corner apex along the bisector. The corner angle is $\psi$, while the metal films have a length $L$ and height $H$, as shown in the figure. The thickness of the metal films forming the reflector is $W$. The planes of the CR are assumed to be made of the same metal as used for the NP in the feeder system.
Figure 1: Geometry employed in the analyses. The feeder is composed by a dipole emitter $\mathbf{p}$ coupled to a nanosphere of diameter $d$. The corner, of angle $\psi$, is assumed to be built from to sheets of the same metal as the NP with dimensions $L\times H$.
A metallic plane typically reflects a substantial portion of an incident wave originating from a large distance, due to the significant mismatch between its optical properties and those of the surrounding dielectric. However, when a noble metal plane is placed near a dipolar source, it can significantly influence the emission rate [21], similar to the behavior observed with emission near NPs. Therefore, analyzing only the radiation pattern of emitted photons is insufficient for designing CR-based NAs. A comprehensive analysis, including the effects on radiative emission and loss rates, has been conducted. The impact of the thickness of the metal films on these properties has also been addressed.
2 Materials and Methods
The numerical calculations were conducted using ANSYS Lumerical FDTD Solutions, release 2024 R2.1, on a HP Z6 G4 Workstation with two Intel Xeon 6226R processors and 512GB of RAM. For the finite-difference time-domain (FDTD) simulations, a variable-size discretization was employed, with a finer grid in the dipole-nanoparticle region. The discretization parameters were adjusted based on a series of convergence analyses, setting the maximum grid step to $\Delta l=1$ nm within this region. To reduce computational costs, the symmetry planes $XZ$ and $YZ$ were utilized for the feeder system, and $YZ$ symmetry was exploited in simulations including the CR. The dipole source emitted a short pulse covering the spectral region between $300$ and $800$ nm. The total duration of the simulations was set to $T=150$ fs. The analyses focused on a target wavelength $\lambda_{Ag}=450$ nm for silver and $\lambda_{Au}=600$ nm for gold, both red-shifted from their respective resonances with similar spectral separations.
For the hybrid nano-antenna in an homogeneous dielectric environment, the enhancement factor of the radiative decay rate, $f_{R}$, due to the Purcell effect [22] can be calculated using a fully classical model [21] as the ratio between the total power collected in the far-field region when the noble metal antenna is present, $P_{rad}$, and the power that would be radiated from a point dipole source in free space, $P_{0}$,
$f_{R}=\dfrac{\gamma_{r}}{\gamma_{r,0}}=\dfrac{P_{rad}}{P_{0}},$ (1)
where $\gamma_{r}$ ($\gamma_{r,0}$) is the radiative emission rate in the presence (absence) of the plasmonic structure. $P_{rad}$ and the radiative emission rate enhancement factor are readily obtained from FDTD computations [23, 9]. For this purpose, the atom-like emitter is modeled using a time-windowed dipole soft source. Instead of setting a given value of the electromagnetic field at a point in space, the field from this type of source adds to the existing electromagnetic fields without causing scattering, therefore permitting the modeling of fluorescence. The value of $P_{0}$ can be computed from its analytical expression [24], but it is also provided by Lumerical via the sourcepower script command.
Similarly, the impact of the losses in the antenna structure, which affect the overall quantum efficiency, can be evaluated from the normalized loss rate factor
$f_{L}=\dfrac{\gamma_{l}}{\gamma_{r,0}}=\dfrac{P_{loss}}{P_{0}}.$ (2)
$P_{loss}$ is computed as $P_{loss}=P-P_{rad}$ where $P$ is the power radiated by the dipole in the NA environment used to modify its the radiation in vacuum, $P_{0}$.
The total photon emission rate is influenced by radiative and non-radiative transitions and losses in the noble metal structures. The emission yield is characterized by quantum efficiency $\eta_{Q}$ of the process: the ratio between the rate of radiative transitions and the total transition rate including radiative $\gamma_{r}$ and non-radiative $\gamma_{nr}$ transitions and the rate of photons lost in the metals $\gamma_{l}$ [21],
$\eta_{Q}=\dfrac{\gamma_{r}}{\gamma_{r}+\gamma_{nr}+\gamma_{l}}.$ (3)
If we assume a highly efficient atom-like emitter, with negligible non-radiative transitions $\gamma_{nr}<<\gamma_{r}$, the quantum efficiency can be approximated as
$\eta_{Q}\simeq\dfrac{\gamma_{r}}{\gamma_{r}+\gamma_{l}}=\dfrac{f_{R}}{f_{R}+f_{L}}$ (4)
where we have divided both numerator and denominator of the above expression by $\gamma_{r,0}$. If we substitute (1) and (2) into (4) the resulting expression coincides with the classical definition of the radiation efficiency of an antenna [15, 16]
$\eta=\dfrac{P_{rad}}{P_{rad}+P_{loss}}$ (5)
The evaluation of the emission properties for the NAs in this study is based on the calculation of $f_{R}$ and $f_{L}$ from the FDTD results to gauge the photon emission rate. Both gold and silver plasmonic NA were considered, with their respective optical constants obtained from [25] and [26].
The external photon collection capabilities of the hybrid system are evaluated using of the on-axis directivity, determined from the far-field radiation pattern as
$D=4\pi\dfrac{K_{0}}{P_{rad}},$ (6)
where $K_{0}=K\left(\theta=\pi/2,\phi=\pi/2\right)$ is the radiation intensity emitted along the $Y$ axis, i.e., the power per unit solid angle in that direction. In Eq. (6), $D$ measures the ratio of the power radiated per unit solid angle along the $Y$-axis to that of a hypothetical isotropic radiator [15, 16].
The antenna gain, defined as
$G=\eta D,$ (7)
combines the emission efficiency $\eta$ and directivity $D$, serving as the key metric for antenna performance in a RF link [15, 16]. This figure of merit is applicable to optical NAs, where it indicates the potential for photon collection from single-photon emitters. We now analyze the values of $D$, $f_{R}$ and $f_{L}$, the three parameters that determine $G$ for silver and gold NAs.
3 Results and Discussion
An initial analysis was performed on the isolated dipole-NP feeder problem. This permits the full evaluation of the impact on the emitter properties of the incorporation of the CR. To this end, a series of FDTD calculations were performed on the geometry displayed on Fig. 1 in the absence of the metal sheets defining the CR.
The resonances in the spectral responses of NPs are responsible for their characteristic colors and their intensities and frequencies depend on the particle size, shape and the dielectric properties of the surrounding medium. Fig. 2 displays the resonances calculated from the emission spectrum of the corresponding NPs coupled to the dipole emitter. For gold NPs surrounded by air, the calculations show a negligible displacement of the plasmon resonance for the range of sizes considered in accordance with previous calculations [27]. Furthermore, The spectral shift of the resonance with the particle size correlate well with the red-shift of the resonance as the particle size is increased observed in measurements performed for colloidal suspensions of silver NPs of sizes varying between $60$ and $80$ nm in [28, 29].
(a) (b)
Refer to caption Refer to caption
Figure 2: Calculated resonance frequencies of silver (a) and gold (b) NPs as a function of their diameter.
Fig. 3 displays the results obtained for the normalized factors $f_{R}$ and $f_{L}$ obtained from FDTD calculations for silver NPs of diameter $d$ coupled to a dipole emitter at a distance $s$ and perpendicular to the spherical NP. Particle sizes between $d=30$ and $80$ nm have been considered and dipole separations ranging from $s=5$ nm to $s=30$ nm. The plots display data collected at the resonance and at the target wavelength of $\lambda_{Ag}=450$ nm. The corresponding results for gold NPs with target wavelength at $\lambda_{Ag}=600$ nm are shown in Fig. 4. In this case, the calculations have been performed for the same range of parameters $d$ and $s$.
(a) (b)
Refer to caption Refer to caption
(c) (d)
Refer to caption Refer to caption
Figure 3: Radiative transition rate enhancement, $f_{R}$, and loss, $f_{L}$, factors of silver NPs calculated for different spherical particle diameter $d$ and separation $s$ from the the dipole emitter. $(a)$ and $(c)$ plots are the values at resonance and $b$ and $d$ at the target wavelength of $\lambda_{Ag}=450$ nm.
(a) (b)
Refer to caption Refer to caption
(c) (d)
Refer to caption Refer to caption
Figure 4: Radiative transition rate enhancement, $f_{R}$, and loss, $f_{L}$, factors of gold NPs calculated for different spherical particle diameter $d$ and separation $s$ from the the dipole emitter. $(a)$ and $(c)$ plots are the values at resonance and $b$ and $d$ at the target wavelength of $\lambda_{Au}=600$ nm.
The calculated values of the radiative transition factor, $f_{R}$, at the target wavelengths generally increase as the particle size grows and the distance between the emitter and the NP diminishes, both for silver and gold NPs. However, the variation of $f_{R}$ with particle size $d$ at resonance differs between silver and gold. For gold NPs, the trend in $f_{R}$ at resonance is similar to that observed at $\lambda_{Au}$, while for silver NPs, $f_{R}$ exhibits a local maximum at a specific $d$ when the separation $s$ is small. In general, $f_{R}$ reaches higher values for silver NPs than for gold, both at resonance and at the target wavelength. Although the relative size of the NPs compared to the wavelength at which $f_{R}$ is measured is larger for silver NPs, the ratio of the relative sizes is smaller than the observed ratios of $f_{R}$ values. On the other hand, when the value of the wavelength is taken in consideration, the corresponding relative distances between the source and the NP are larger for silver NPs.
Regarding photon loss due to absorption at the NP, the value of $f_{L}$ increases significantly as the distance between the dipole emitter and the NP decreases. Conversely, the NP size $d$ shows a lesser effect on $f_{L}$. Overall, the loss factors for silver NPs are smaller, though still comparable, to those for gold NPs.
As discussed in Section 2, achieving efficient emission requires a good balance between relatively low $f_{L}$ values and sufficiently high $f_{R}$ values. Based on the analysis of the coupled dipole-nanoparticle system, a particle size of $d=70$ nm with a separation of $s=15$ nm from the dipole was chosen for the analysis of the full structure incorporating the CR. For these parameters, at their respective target wavelengths, the results show $f_{L}=5.4$ and $f_{R}=8.3$ for silver nanoparticles, and $f_{L}=6.1$ and $f_{R}=4.9$ for gold nanoparticles. This indicates a higher radiative emission rate enhancement with lower losses for silver compared to gold.
Under these conditions, $f_{R}>f_{L}$ for silver NPs whereas $f_{R}<f_{L}$ for gold NPs. The corresponding values of the emission gain figure of merit, $G$, as defined in (7), with a directivity value of $D=1.5$ for an electrically short dipole [15, 16], is $G_{Ag}=0.9$ for silver and $G_{Au}=0.7$ for gold NPs.
The radiation properties of CRs in RF antennas were systematically studied in [30] through extensive experimental measurements of the gain achieved with different corner configurations using a resonant half-wave dipole feeder. As the distance from the corner apex to the feeder increases, a series of positions corresponding to local maxima of the antenna gain are identified. The first three positions were thoroughly analyzed in [30].
Whereas the second and third positions are sharply defined, the first position corresponds to a broader region. This is due to a trade-off between the increase in directivity obtained by moving the feeder closer to the apex and the associated rise in ohmic losses in the conductors. In the first position, the corner angle for maximum gain depends on the width and length of the reflecting planes: as the planes become larger, the optimal angle decreases and the maximum gain increases. Conversely, fixed optimal values of the corner angle are observed for the second and third positions when the planes are sufficiently large. For the second position, a maximum gain occurs at $\psi=65^{\circ}$ when the width and length of the corner planes exceed about one wavelength. At the third position, an optimal corner angle is found at $\psi=90^{\circ}$ when the reflecting planes are larger than half a wavelength.
In the following analyses, we select a corner angle of $\psi=65^{\circ}$ and corner plane dimensions of $L=1.5$ $\mu$m and $H=1$ $\mu$m. We examine the directivity $D$ along the corner axis and investigate how the presence of the CR influences the values of $f_{R}$ and $f_{L}$ at the target wavelengths for silver and gold NAs.
(a) (b)
Refer to caption Refer to caption
Figure 5: Values of the directivity as a function of the feeder position $S$, normalized to the respective target wavelength, for four different reflector film thicknesses. (a) Results for the silver CR, with a target wavelength of $\lambda_{Ag}=450$ nm. (b) Results for the gold CR, with a target wavelength of $\lambda_{Au}=600$ nm.
Fig. 5 presents the on-axis directivities calculated for silver (a) and gold (b) CRs, with film thicknesses ranging from $W=10$ nm to $80$ nm and feeder positions between $S=0.2$ $\mu$m and $S=1$ $\mu$m. For the presentation of the results, the values of $S$ have been rescaled with the respective target wavelengths for the gold and silver reflectors at which the directivities are evaluated.
The figure reveals similar patterns in the variation of directivity with the normalized separation $S/\lambda_{M}$ and the film thickness $W$, where $M$ denotes the corresponding noble metal. The higher peak directivities obtained for silver than for gold can be attributed mainly to the larger relative values of the corner plane length $L/\lambda_{Ag}>L/\lambda_{Au}$.
The metal film thickness also plays a significant role in the radiation properties of the NA. As the thickness increases, the peak directivity increases, and the variation with $S/\lambda_{M}$ follows a more regular oscillatory pattern with local maxima and minima as in the case of RF antennas [30]. The irregularities in variations of directivity at smaller $W$ are less pronounced for gold mirrors compared to silver. Additionally, the positions of maxima and minima shift to larger values of $S/\lambda_{M}$ as $W$ increases. At the largest thickness of $W=80$ nm, there is directivity null precisely at $S=\lambda_{M}$.
A better understanding of the variations observed in the curves of $D$ can be achieved through a detailed analysis of the radiation patterns in each case. Such plots are presented for silver CRs with $W=10$ nm, $W=40$ nm and $W=80$ nm ,respectively, in figures 6, 7 and 8.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 6: Radiation patterns of silver CR with $W=10$ nm for values of $S$ between $S=0.2$ and $S=0.75$.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 7: Radiation patterns of silver CR with $W=40$ nm for values of $S$ between $S=0.2$ and $S=0.75$.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 8: Radiation patterns of silver CR with $W=80$ nm for values of $S$ between $S=0.2$ and $S=0.75$.
One of the main differences observed at small film thickness ($W=10$ nm) is the presence of relevant amplitude side lobes at the lower values of $S$. This phenomenon leads to reduced directivity values, despite the main lobe exhibiting an narrower beamwidth at $W=10$ nm. Conversely, the patterns seen at $W=40$ nm and $W=80$ nm are qualitatively similar, indicating a convergence as the thickness $W$ increases.
For thicker sheet CRs, the gradual narrowing (or broadening) of the on-axis lobe contributes to regions of increasing (or decreasing) directivity values. The zero observed in the directivity curves arises from the splitting of the main lobe at $S=\lambda_{M}$ that produces a radiation null along the $Y$ axis. In contrast, for $W=10$, the local minima in the directivity plots correlate with the increasing relative weight of the side lobes in the radiation pattern. As $S$ increases, separation of the main lobe is also noted; however, it occurs along the horizontal plane, unlike the vertical splitting seen at $W=40$ nm and $W=80$ nm. As $S$ continues to grow, additional lobe bifurcations become apparent, leading to the emergence of a fractal fractal structure in the radiation pattern for very large values of $S$ in extensive CRs, as analyzed in [18].
The sequences of radiation diagrams presented in Figs. 6, 7, and 8 are very similar for the corresponding ones of gold CRs when the values of $S$ are normalized by the respective operation wavelength $\lambda_{M}$.
We now examine the impact of the corner reflector on the losses and the emission rate of the dipole-NP system. Fig. 9 displays $f_{R}$, and the normalized loss factor $f_{L}$ of NAs incorporating silver CRs of four different thicknesses, $W$, as a function of the normalized distance, $S/\lambda_{Ag}$, between the silver NP and the corner apex. The corresponding results for gold NAs are shown in Fig. 10. In all the figures, the horizontal lines pinpoint the reference values for the NPs without the CRs.
(a) (b)
Refer to caption Refer to caption
(c) (d)
Refer to caption Refer to caption
Figure 9: Radiative transition rate enhancement $f_{R}$ and normalized loss $f_{L}$ factors of a silver CR as a function of the normalized feeder position $S/\lambda_{M}$ for four different reflector film thicknesses: (a) $W=10$ nm, (b) $W=20$ nm, (c) $W=40$ nm and (d) $W=80$ nm. The horizontal lines describe the corresponding reference levels without the CR.
(a) (b)
Refer to caption Refer to caption
(c) (d)
Refer to caption Refer to caption
Figure 10: Radiative transition rate enhancement $f_{R}$ and normalized loss $f_{L}$ factors of a gold CR as a function of the feeder position $S/\lambda_{M}$ for four different reflector film thicknesses: (a) $W=10$ nm, (b) $W=20$ nm, (c) $W=40$ nm and (d) $W=80$ nm.
For silver CRs, the normalized loss factors display moderate deviations from the reference values except for very small values of $S/\lambda_{Ag}$, where a steep increase of the losses occurs due to stronger interactions with the nearby CR metal films when the thickness, $W$, is sufficiently large. In contrast, for gold CRs, Fig. 10 indicates larger fluctuations in the loss factor around the reference levels.
The calculated radiative emission rate enhancement factors also exhibit a pronounced effect due to the interaction with the corner sheets when $W$ is sufficiently large. This interaction significantly impacts both silver and gold NAs, and the associated $S/\lambda_{M}$ range is larger than that observed for $f_{L}$ in silver NRs, increasing with $W$. The variation with $S/\lambda_{M}$ follows a distinct pattern: $f_{R}$ reaches a maximum near the first peak of directivity, then drops to a minimum around the directivity null at $S/\lambda_{M}=1$, before rising again near the second peak.
For silver CRs with $W=80$ nm, we find that near the first directivity maximum at $S=0.30$ $\mu$m ($S/\lambda_{Ag}=0.67$), the values are $D=20.1$, $f_{R}=14.5$, and $f_{L}=6.7$. The second maximum yields a significantly higher directivity of $D=40.4$ at $S=0.55$ $\mu$m ($S/\lambda_{Ag}=1.22$), with $f_{R}=6.0$ and $f_{L}=5.9$. Thus, the CR provides very large improvements of the directivity when compared to that of a electric dipole emitter, $D=1.5$.
Considering the reference values for the NP emitter, $f_{R}=8.3$ and $f_{L}=5.4$, the first maximum offers a significant increase in both directivity and radiative emission rate enhancement factor, with only a modest rise in losses. At the second maximum, the directivity is much larger, though there is a slight reduction in $f_{R}$ in relation to the reference. The loss factor, however, remains close to that of the emitter without the CR. The gain figure of merit is $G=13.8$ and $G=20.4$ for the first and second maxima, respectively, largely exceeding the value $G=0.9$ obtained for the NA without the CR.
For the gold NA, the reference values $f_{R}=4.9$ and $f_{L}=6.1$. With a $W=80$ nm CR NA, at the first maximum, located at $S=0.35$ $\mu$m ($S/\lambda_{Au}$=0.58), the directivity is $D=18.0$, with $f_{R}=8.2$ and $f_{L}=7.0$. The second maximum occurs at $S=0.75$ $\mu$m ($S/\lambda_{Au}=0.58$), where $D=29.0$, $f_{R}=4.4$ and $f_{L}=6.7$.
For gold CRs, the first maximum provides a high directivity and an improvement of $f_{R}$. At the second maximum, while the directivity is even greater, $f_{R}$ is close to the reference value. In both cases, the losses are similar to those of the gold NP without the CR. The gain figure of merit is $G=8.0$ and $G=11.5$ for the first and second maxima, respectively, significantly exceeding the value $G=0.7$ obtained for the NA without the CR but significantly lower than the values obtained for silver CR NAs. Besides the impact of the better figures for $f_{R}$ and $f_{L}$ in the case of silver NAs, the lower values of $G$ obtained in the case of gold CR NAs is also due to the fact that $L/\lambda_{M}$ is smaller in the case of gold.
(a) (b)
Refer to caption Refer to caption
Figure 11: Values obtained of the directivity as a function of the feeder position $S$ normalized to the corresponding target wavelength for three different reflector film thicknesses when the length of the CR planes is increased to $L=3$ $\mu$m. (a) Results for silver CR. (b) Results for gold CR.
Fig. 11 shows the directivities calculated for silver and gold CR NAs when the corner length is increased to $L=3$ $\mu$m for three values of $W$ ($10$ nm, $40$ nm and $80$ nm). For this $L$, the peak directivities are essentially the same for silver and gold CRs, even though $L/\lambda_{M}$ is larger for silver. This indicates a saturating effect in the increase of $D$ with $L$.
In the case of gold, the relative positions of the maxima and minima as a function of $S/\lambda_{M}$ closely match those found for $L=1.5$ $\mu$m, with a null occurring at $S=\lambda_{M}$ for $W=80$ nm. In contrast, for silver, significant shifts in the directivity maxima and minima as a function of $S/\lambda_{M}$ are observed , and these shifts are highly dependent on the value of $W$.
(a) (b)
Refer to caption Refer to caption
(c) (d)
Refer to caption Refer to caption
(e) (f)
Refer to caption Refer to caption
Figure 12: Radiative transition rate enhancement $f_{R}$ and normalized loss $f_{L}$ factors of a gold CR with $L=3$ $\mu$m as a function of the feeder position $S/\lambda_{M}$ for three different reflector film thicknesses: (a) silver CR $W=10$ nm, (b) gold CR $W=10$ nm, (c) silver CR $W=40$ nm, (d) gold CR $W=40$ nm, (d) silver CR $W=80$ nm, (d) gold CR $W=80$ nm.
The normalized radiative emission and loss rates for $L=3$ $\mu$m CRs are displayed in Fig. 12 for the three values of $W$ considered. The results closely resemble those obtained for $L=1.5$ $\mu$m, indicating that the impact of the CR on the loss and emission properties is due to the interactions occurring in a localized region near the apex of the CR.
Fig. 13 shows the gain, $G$, of silver and gold CR NAs as a function of $S/\lambda_{M}$. The gain provides a comprehensive measure of on-axis photon emission, incorporating all three key parameters: $f_{R}$, $f_{L}$, and $D$. For gold CR NAs, the maximum gain occurs at a metal thickness of $W=80$ nm. The first gain peak, with a value of $G=11.8$, is found at $S/\lambda_{Au}=0.67$ and $S/\lambda_{Au}=0.75$. The second peak, $G=21.8$, is located at $S/\lambda_{M}=1.33$.
In the case of silver, the aforementioned shift of the directivity maxima with respect to $S/\lambda_{Ag}$ results in the highest directivity for $W=40$ nm rather than $W=80$ nm. This shift aligns the directivity peak with a higher $f_{R}$ value at $W=40$ nm, leading to the maximum gain being achieved at this thickness. The first gain peak for silver is $G=12.3$ at $S/\lambda_{Ag}=0.5$ and the second is $G=30.0$ at $S/\lambda_{Ag}=1.25$.
(a) (b)
Refer to caption Refer to caption
Figure 13: Nanoantenna gain as a function of the feeder position $S$ normalized to the corresponding target wavelength for three different reflector film thicknesses when the length of the CR planes is increased to $L=3$ $\mu$m. (a) Results for silver CR. (b) Results for gold CR.
4 Conclusions
The analysis of dipole-NP interactions, combined with CR integration, reveals important insights into the enhancement of radiative properties for silver and gold NPs. The fundamental behavior of NP interactions is crucial for understanding the modifications introduced by the incorporation of the CR in the system. Initially, isolated dipole-NP calculations without the CR showed trends consistent with previous theoretical and experimental studies, particularly regarding the variation in resonance wavelength, the radiative emission rate enhancement factor, and loss factor with silver and gold particle size and dipole emitter separation. This preliminary study established a set of parameters for the subsequent analysis of the CR. The results also suggest that silver NPs are more effective than gold NPs in enhancing radiative emission while maintaining lower loss factors.
Integrating CRs into silver and gold NA system leads to substantial improvements in directivity. The variation in directivity with respect to the normalized distance $S/\lambda_{M}$ reveals that thicker metal films generally lead to more regular oscillatory patterns in directivity. At the first directivity maxima, significant gains in the radiative enhancement factor, $f_{R}$, are also achieved with minimal increases in the loss factor, $f_{L}$, leading to high gain figures of merit, $G$. At the second directivity maximum, the CR mainly enhances directivity, with the highest values of $D$ corresponding to the global maxima of $G$.
Extending the corner length to $L=3$ $\mu$m demonstrates a saturation effect in the enhancement of directivity for both materials, with similar peak values achieved for gold and silver CRs despite differences in $L/\lambda_{M}$. Meanwhile, the patterns for the variation of $f_{R}$ and $f_{L}$ closely resemble those found at $L=1.5$ $\mu$m. These findings highlight the localized nature of CR-induced enhancements near the apex. However, silver CRs exhibit larger shifts in the positions of directivity maxima and minima with changes in $W$. This results in the maximum gain for silver CR NAs at a metal thickness $W=40$ nm, rather than $W=80$ nm, due to the alignment of the directivity maximum with a higher value of $f_{R}$.
Overall, the results indicate that silver CR NAs outperform gold CR NAs in radiative enhancement and gain, offering more efficient light emission and directivity control for nanoscale optical applications, although different spectral windows are targeted by the choice of NA metal.
The reflector angle selected for the analysis aligns with a typical value used in RF applications, corresponding to optimization for operation at the more favorable second gain maximum. The analysis reveals a strong metal-feeder interaction near the corner apex, suggesting the potential benefits of further exploring corner geometries to optimize the first gain maximum, which will be addressed in future works.
Acknowledgement
This work was funded by the Spanish Ministerio de Ciencia e Innovación (MCIN) under grant PID2020-119418GB-I00, by the European Union NextGeneration EU under grant PRTRC17.I1, and by the Consejería de Educación, Junta deCastilla y León, through QCAYLE project, and grant VA184P24.